Equivalent Units of Production and Related Costs The charges to Work in Process-Baking Department for a period as well as information concerning production are as follows. The Baking Department uses the average cost method, and all direct materials are placed in process during production Work in Process-Baking Department Bal., 3,300 units, 60% completed Direct materials, 24,100 units Direct labor Factory overhead Bal., 3,700 units, 40% completed If required, round cost per equivalent unit answer to two decimal places. 8,250 To Finished Goods, 23,700 units 87,965 42,858 22,079 a. Determine the number of whole units to be accounted for and to be assigned costs units b. Determine the number of equivalent units of production units c. Determine the cost per equivalent unit. per equivalent unit d. Determine the cost of the units transferred to Finished Goods. e. Determine the cost of units in ending Work in Process.
Answer #1

a. The number of whole units to be accounted for and to be assigned costs = 23,700 + 3,700 = 27,400 units

b. The number of equivalents units of productions

= 23,700 + (3,700 x 40%) = 25,180 units

c. The cost per equivalent unit

= (8,250 + 87,965 + 42,858 + 22,079) ÷ 25,180 = $6.4 per equivalent unit

d. The cost of the units transferred to Finished Goods

= 23,700 x 6.4 = $151,680

e. The cost of units in ending Work in Process

= (3,700 x 40%) x 6.4 = $9,472